A container encloses 31.6 g of CO2(g). The partial pressure of CO2 is 2.61 atm and the volume of the container is 34.9 L. What is the average, or root mean square, speed (in m/s) of the CO2 molecules in this container?
1
Expert's answer
2020-07-16T06:59:59-0400

"v_{rms}=\sqrt{\frac{3RT}{M}}=\sqrt{\frac{3PV}{nM}}=\sqrt{\frac{3\times2.61\times34.9}{31.6\times10^{-3}}}=93m/s"